OpenStudy (anonymous):

evaluate the integral: ∫cos²x on (0, π/4)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Hints: \[\cos^2x=(1+\cos2x)/2\] this is the one of the Power-Reducing Formulas. then you integrate it.

OpenStudy (anonymous):

woah my teacher never taught me that formula.. i don't think she would want me to do it in a way she hasn't taught

OpenStudy (anonymous):

Then I can't help you further than than :(

OpenStudy (anonymous):

darnit :( is it a fairly easy formula to understand?

OpenStudy (anonymous):

do you know this?\[\cos(2x)=2\cos^2x-1\]

OpenStudy (anonymous):

no i haven't learned that

OpenStudy (anonymous):

then might be a little hard to explain...

OpenStudy (anonymous):

sorry i have to go to bed but I will get to class early tomorrow to ask my teacher!

OpenStudy (anonymous):

you can go to this website to have a see http://jwilson.coe.uga.edu/emt668/EMAT6680.2003.fall/Bismarck/6690%20folder/trigproofs/triproofs.html

OpenStudy (anonymous):

ok~good night~